OpenStudy (anonymous):

What is the solution to -5+2x^2=-6x

OpenStudy (anonymous):

move -6x to the left-hand right then use the quadratic for formula to solve for roots.

OpenStudy (anonymous):

2x^2 +6x -5 = 0 factor using techniques or use qudratice formula. (x - 6)(x+ 1) and your solutions are x = 6 and x - 1

OpenStudy (phebe):

2 x (x+3) = 5